OpenStudy (anonymous):

64, 16, 4, 1, whats the common ratio?

OpenStudy (goformit100):

First of all @rosiitas : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

\(64\times r=16\) so \(r=\frac{16}{64}\)

OpenStudy (anonymous):

it is always the next term over the previous one \[\frac{16}{64}=\frac{4}{16}=\frac{1}{4}=...\]

OpenStudy (anonymous):

Welcome 2 os